OpenStudy (anonymous):

If P dollars are deposited at an interest rate r and compounded n times, the future value An can be found by the formula An=P(1+r)n. find the rate of interest if a principal amount of $4500 grows to $5644.80 in 2 years if interest is compounded annually

OpenStudy (anonymous):

Just need some help getting started on this one.

OpenStudy (anonymous):

Compounded annually means that your n = (2)(12) and your r = interest/12. So solve for r and then find I by multiplying by 12. \[5644.80 = 4500(1 + r)^{24}\] solve for r

OpenStudy (anonymous):

divide 5644.80 by 4500 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok...the ^24 is confusing me. so far: (1+r)^24=1.2544

OpenStudy (anonymous):

yes take the 24th root of 1.2544

OpenStudy (anonymous):

the just solve for the positive result of 1-r

OpenStudy (anonymous):

I mean 1+r

OpenStudy (anonymous):

Well, I did something wrong...did not get right answer.

OpenStudy (anonymous):

what did you do?

OpenStudy (anonymous):

1.2544^24=230.39

OpenStudy (anonymous):

no that is wrong

OpenStudy (anonymous):

yep, it's wrong all right..

OpenStudy (anonymous):

\[24\sqrt{1.2544}\]

OpenStudy (anonymous):

you should get 1.009488793

OpenStudy (anonymous):

or 10%?

OpenStudy (anonymous):

rounded

OpenStudy (anonymous):

no that is not done yet

OpenStudy (anonymous):

subtract 1 and multiply it by 12.

OpenStudy (anonymous):

of course not, can't be that simple :)

OpenStudy (anonymous):

11.38 % about is that an option... think more clearly.. there is no way the interest can be 230.39 lol

OpenStudy (anonymous):

That's my problem, math and I do not speak the same language! lol

OpenStudy (anonymous):

that's fine.. practice makes perfect.

OpenStudy (anonymous):

ok, that does make way more sense.

OpenStudy (anonymous):

Thank you, now off to practice some more of these great word problems!

OpenStudy (anonymous):

have fun ;) cya yw